A seven digit number is formed by using digits 1,2,3,4,5,6,7 without repetition . The probability that the number so formed is divisible by 4 is ?

Grade:Upto college level

1 Answers

SHAIK AASIF AHAMED
9 years ago
Hello student,
Seven digit number is formed without repetition so no of ways
=7*6*5*4*3*2*1=5040
for any number to b divisible by 4, the last two digit must
b divisible by 4. Such 2 digit combinations possible are.....
12,16,24,32,36,52,56,64,72,76.
At last two places, any one combination out of these 10 is
possible. For first five digits in the 7 digit number, the
possible combinations are 5*4*3*2*1=120
So there r 120*10=1200 numbers divisible by 4.
So required probability=1200/5040=0.238
Thanks and Regards
Shaik Aasif
askIITians faculty